Let f (x) = sin(ax + b) and let g (x) = cos(ax + b) where a and b are constants. Gues a formular for f*(x) and g"(x) for general n (positive integer )
1
Expert's answer
2020-06-16T19:16:03-0400

Given "f (x) = sin(ax + b), g(x) = cos(ax+b)"

"\implies f'(x) = a \ cos(ax+b), g'(x) = -a \ sin(ax+b)"

"\implies f''(x) = -a^2 sin(ax+b), g''(x)=-a^2cos(ax+b)"

"\implies f'''(x) = -a^3 cos(ax+b), g'''(x) = a^3 sin(ax+b) \\
\implies f^{iv}(x) = a^4 sin(ax+b), g^{iv}(x) = a^4 cos(ax+b)"

Hence In general,

"f^n(x) = \begin{cases} (-1)^{\frac{n}{2}} a^n sin(ax+b) : n \ is \ even \\
(-1)^{\frac{n-1}{2}} a^n cos(ax+b) : n \ is \ odd
\end{cases}"

and "g^n(x) = \begin{cases} (-1)^{\frac{n}{2}} a^n cos(ax+b) : n \ is \ even \\
(-1)^{\frac{n+1}{2}} a^n sin(ax+b) : n \ is \ odd
\end{cases}"